# RSEND: Retinex-based Squeeze and Excitation Network with Dark Region Detection for Efficient Low Light Image Enhancement

### v1

@@ 3.5.4 Perceptual Loss ($\mathcal{L}_{vgg}$) [13]
$\Phi_{l}$ denotes the feature maps from the $l^{th}$ layer of VGG-19, $M_{l}$ is the number of elements in each layer, $I_{gt}$ is the ground truth image, and $L$ is the total number of layers used.
3.5.5 Charbonnier Loss ($\mathcal{L}_{char}$) [15]
Finally, we incorporate the Charbonnier loss, a smooth approximation of the L1 loss, to measure the pixel-wise difference between the enhanced image and the ground truth:
$\mathcal{L}_{char}=\sqrt{\|\hat{I}-I_{gt}\|^{2}+\epsilon^{2}}$ (10)
where $\epsilon$ is a small constant ensuring numerical stability. In our case, we set $\epsilon$ as $31e-$ 3.
[…]
The total loss function $\mathcal{L}_{total}$ is a weighted combination of the Spatial Consistency Loss, Exposure Control Loss, Color Constancy Loss, VGG Loss, and Charbonnier Loss. It is computed as:
$\mathcal{L}_{total}=\lambda_{spa}\cdot\mathcal{L}_{spa}+\lambda_{exp}\cdot\mathcal{L}_{exp}+\lambda_{col}\cdot\mathcal{L}_{col}+\lambda_{VGG}\cdot\mathcal{L}_{VGG}+\lambda_{char}\cdot\mathcal{L}_{char},$ (11)
where $\lambda_{spa}$, $\lambda_{exp}$, $\lambda_{col}$, $\lambda_{VGG}$, and $\lambda_{char}$ are the weights of the losses.
4 Experiment
4.1 Datasets and Implementation details
[…]
@@ 4.1 Datasets and Implementation details
SID [3] 13.73 7.76 14.35     0.436 13.24     0.442 15.04     0.610
Zero-DCE [8] 4.00 0.08 14.86     0.667 18.06     0.680 17.76     0.838
RF [14] 46.23 21.54 15.23     0.452 14.05     0.458 15.97     0.632
DeepLPF [18] 5.86 1.77 15.28     0.473 14.10     0.480 16.02     0.587
UFormer [20] 12.00 5.29 16.36     0.771 18.82     0.771 19.66     0.871
RetinexNet [21] 587.47 0.84 16.77     0.560 15.47     0.567 17.15     0.798
EnGAN [12] 61.01 114.35 17.48     0.650 18.23     0.617 16.57     0.734
RUAS [16] 0.83 0.003 18.23     0.720 18.37     0.723 16.55     0.652
FIDE [22] 28.51 8.62 18.27     0.665 16.85     0.678 15.20     0.612
DRBN [24] 48.61 5.27 20.15     0.830 20.29     0.831 23.22     0.927
KinD [29] 34.99 8.02 20.86     0.790 14.74     0.641 13.29     0.578
Restormer [26] 144.25 26.15 22.43     0.823 19.94     0.827 21.41     0.830
SNR-Net [23] 26.35 4.01 24.61     0.842 21.48     0.849 24.14     0.928
Retinexformer [2] 15.57 1.61 25.16     0.845 22.80     0.840 25.67     0.930
Ours 17.99 0.41 23.67     0.876 24.49     0.895 23.66     0.912
Table 1: Quantitative comparisons on LOL (v1 [21] and v2 [25]) datasets
4.2 Quantitative Results
We compare RSEND with a wide range of state-of-the-art low-light image enhancement networks. Our result significantly outperforms SOTA methods on 3 datasets while requiring much less computational and memory cost shown in Table 1.
When compared with the best CNN-based model DRBN [24], our model achieves 3.54, 4.2, and 0.44 dB improvements on LOL-v1, LOL-v2-real, and LOL-v2-synthetic datasets. Besides, our model only consumes 7.8% (0.41/5.27) parameters and 37 % (17.99/48.61) FLOPS, which is significantly smaller than many of the other high-performing models, highlighting the efficiency of our architecture. When compared with the SOTA transformer-based model Retinexformer [2], our model’s performance in LOL-v2-real dataset yields improvement for 1.69 dB while only consuming 25% (0.41/1.61) parameters FLOPS. Apart from PSNR, our model achieves SSIM scores of 0.876, 0.895, and 0.912 on each dataset, which indicates that our model not only accurately restores brightness levels but also maintains structural integrity and texture details that are crucial for perceptual quality.
4.3 Visual and Perceptual Comparisons
Fig. 3. shows the visual comparisons of the low-light image (left), the other model’s performance (middle), and our RSEND’s performance (right). We can see our model either makes the image lighter or detects more details, showing the effectiveness of the model in different datasets.
@@ 4.3 Visual and Perceptual Comparisons
Figure 3: Visual comparisons with Retinexformer [2], DC-Net [30], EnGAN [12], MIR-Net [27], RetinexNet [21], our RSEND performs better
4.4 Ablation Study
We perform several ablation studies to demonstrate the effectiveness of each part of our network on the LOL-v2-synthetic dataset for its stable convergence. The examples are presented in Fig. 4.
@@ 4.4 Ablation Study
(b) no residual
(c) no SEB
(d) no denoise
(e) Full RSEND
Figure 4: Ablation study of the effect of each component
4.4.1 Residual learning alike algorithm
We conduct an ablation to study the effectiveness of adding the original image in the end. When we just perform element-wise multiplication and denoising, our model yields 23.06 dB in PSNR and 0.902 in SSIM. However, if we add the original image after multiplication, the PSNR and SSIM values are 23.66 dB and 0.912. The difference in Fig. 4(b) and Fig. 4(e) shows that with this residual learning alike algorithm, the model achieves an improvement of 0.6 dB in PSNR and 0.01 in SSIM.
4.4.2 SEBlock performance
We conduct an ablation study to evaluate the necessity of adding SEBlock. In Fig. 4(c) and Fig. 4(e), we can clearly see the difference. The output without SEBlock is not well lighted and we can not find visually pleasing details. By adding SEBlock, the PSNR and SSIM values gain an improvement of 2.81 dB and 0.037, suggesting the necessity of using SEBlock.
4.4.3 Denoise phase in the end
We conduct an ablation to study the effect of Denoising after using Retinex theory to get the final image. As we can see in Fig. 4(d) and Fig. 4(e), even though the output without the denoising phase preserves relatively visually pleasing results, we can still find in some darker areas the details are missing. We get an improvement of 1.9 dB in PSNR and 0.02 in SSIM, which proves the layer’s efficacy in mitigating noise and preserving detail.
5 Conclusion
@@
We propose an efficient and accurate CNN-based deep learning network for low-light image enhancement and it can be trained end to end with paired image. We start with Retinex theory, dividing the image into reflectance and illumination first. In the process of enhancing, we add our illumination constraints. First, we make the model understand which parts are darker and require more attention by introducing the dark region detection module. After enhancing, we refine the output, which aims to fine-tune the details. After element-wise multiplication of reflectance and illumination, we add the original low-light image in the end, which serves as residual learning to maintain high similarity. Then we denoise the output image to ensure the final result is not only well-illuminated but also visually pleasing. For all the steps mentioned above, we utilize Squeeze and Excitation Network to better capture the details. The quantitative and qualitative experiments show that our RSEND outperforms all the CNN-based models and yields results that are close to the transformer-based models while using very few parameters.

### v2

@@ IV-A Datasets and Implementation details
TABLE I: Quantitative comparisons on LOL (v1 [2] and v2 [22]) datasets
@@
SID
PSNR     SSIM
SID [1] 13.73 7.76 14.35     0.436 13.24     0.442 15.04     0.610 16.97     0.591
Zero-DCE [16] 4.00 0.08 14.86     0.667 18.06     0.680 17.76     0.838 13.68     0.49
RF [23] 46.23 21.54 15.23     0.452 14.05     0.458 15.97     0.632 16.44     0.596
DeepLPF [24] 5.86 1.77 15.28     0.473 14.10     0.480 16.02     0.587 18.07     0.600
UFormer [25] 12.00 5.29 16.36     0.771 18.82     0.771 19.66     0.871 18.54     0.577
RetinexNet [2] 587.47 0.84 16.77     0.560 15.47     0.567 17.15     0.798 16.48     0.578
EnGAN [3] 61.01 114.35 17.48     0.650 18.23     0.617 16.57     0.734 17.23     0.543
RUAS [9] 0.83 0.003 18.23     0.720 18.37     0.723 16.55     0.652 18.44     0.581
FIDE [26] 28.51 8.62 18.27     0.665 16.85     0.678 15.20     0.612 18.34     0.578
DRBN [4] 48.61 5.27 20.15     0.830 20.29     0.831 23.22     0.927 19.02     0.577
KinD [5] 34.99 8.02 20.86     0.790 14.74     0.641 13.29     0.578 18.02     0.583
Restormer [27] 144.25 26.15 22.43     0.823 19.94     0.827 21.41     0.830 22.27     0.649
SNR-Net [28] 26.35 4.01 24.61     0.842 21.48     0.849 24.14     0.928 22.87     0.625
Retinexformer [8] 15.57 1.61 25.16     0.845 22.80     0.840 25.67     0.930 24.44     0.680
RSEND (ours) 17.99 0.41 24.18     0.860 23.92     0.867 24.91     0.912 22.40     0.775
IV-B Quantitative Results
We compare RSEND with a wide range of state-of-the-art low-light image enhancement networks. Our result significantly outperforms CNN-based SOTA methods on these four datasets, while requiring much less computational and memory cost; the comparison is shown in Table I.
Compared with the best CNN-based model DRBN [4], our model achieves 4.03, 3.63, 1.69, and 3.38 dB improvements on the LOL-v1, LOL-v2-real, LOL-v2-synthetic, and SID datasets. In addition, our model only consumes 7.8% (0.41/5.27) parameters and 37 % (17.99/48.61) FLOPS, which is significantly smaller than many of the other high-performing models, highlighting the efficiency of our architecture. Compared with the SOTA transformer-based model Retinexformer [8], the performance of our model in the LOL-v2-real dataset yields an improvement of 1.12 dB while only consuming 25% (0.41/1.61) of the parameters. Apart from PSNR, our model achieves SSIM scores of 0.860, 0.867, 0.912, and 0.775 in each dataset, indicating that our model not only accurately restores brightness levels, but also maintains structural integrity and texture details that are crucial for perceptual quality.
IV-C Visual and Perceptual Comparisons
Fig. 4. shows the visual comparisons of the low-light image (left), the other model’s performance (middle), and our RSEND’s performance (right). We can see that our model either makes the image lighter or detects more details, showing the effectiveness of the model in different datasets.
@@ IV-C Visual and Perceptual Comparisons
(p) Input
(q) ZeroDCE
(r) RSEND
Figure 4: Visual comparisons with Retinexformer [8], DC-Net [29], EnGAN [3], MIR-Net [30], RetinexNet [2], Zero-DCE [16] our RSEND performs better.
IV-D Ablation Study
We perform several ablation studies to demonstrate the effectiveness of each part of our network on the LOL-v2-synthetic dataset for its stable convergence. The examples are presented in Fig. 5 and Table II.
TABLE II: Ablation Study of RSEND Components. The table shows the impact of removing each component on PSNR and SSIM performance. ✓indicates the presence of the component, while empty denotes its removal.
Method Variation SEBlock Dark Region Detection Residual Refinement Denoising PSNR (dB) SSIM
Baseline RSEND 24.91 0.912
w/o SEBlock 20.85 0.875
w/o Dark Region Detection 21.75 0.890
w/o Residual 23.06 0.902
w/o Refinement 22.13 0.882
w/o Denoising 21.90 0.880
@@ IV-D Ablation Study
(b) w/o residual
(c) w/o SEB
(d) w/o denoise
(e) RSEND
(f) Input
(g) w/o dark
(h) w/o refine
(i) RSEND
Figure 5: Ablation Study of the effect of each model component
(a) Input
(b) $\mathcal{L}_{col+spa+exp}$
(c) $\mathcal{L}_{Charbonnier}$
(d) $\mathcal{L}_{comb}$
(e) RSEND
Figure 6: Ablation Study of the effect of different loss functions
IV-D1 Each Component of the Pipeline
We conduct ablations to study the effectiveness of each component of the model pipeline. The first is adding the original image at the end. When we just perform element-wise multiplication and denoising, our model yields 23.06 dB in PSNR and 0.902 in SSIM. However, if we add the original image after multiplication, the PSNR and SSIM values are 24.91 dB and 0.912. The difference in Fig. 5(b) and Fig. 5(e) shows that with this residual learning-alike algorithm, the model achieves an improvement of 1.85 dB in PSNR and 0.01 in SSIM. The second is the necessity of adding SEBlock. In Fig. 5(c) and Fig. 5(e), we can clearly see the difference. The output without SEBlock is not well lighted and we can not find visually pleasing details. By adding SEBlock, the PSNR and SSIM values gain an improvement of 4.06 dB and 0.037, suggesting the necessity of using SEBlock. The third is the effect of Denoising after using Retinex theory to get the final image. As we can see in Fig. 5(d) and Fig. 5(e), even though the output without the denoising phase preserves relatively visually pleasing results, we can still find in some darker areas the details are missing. We get an improvement of 3.15 dB in PSNR and 0.02 in SSIM, which proves the layer’s efficacy in mitigating noise and preserving detail. For the rest of the pipelines, as shown in Fig. 5(g), removing the Dark Region Attention Module results in a significant loss of detail in the darker areas of the image. Even though the image becomes brighter, the color and exposure look unnatural overall, and some parts of the clouds are even black. This demonstrates the module’s effectiveness in enhancing visibility in underexposed regions. As for the image in Fig. 5(h), without the Refinement Layer, it shows noticeable artifacts and less smooth transitions in lighting, as reflected by a PSNR of 22.13 dB and an SSIM of 0.882, which demonstrates its role in reducing noise and enhancing detail for the enhanced output. Our full RSEND in Fig. 5(i) exhibits balanced lighting, enhanced detail, and color accuracy. This result is achieved by integrating all model components, showcasing the synergistic effect that our architectural design aims to accomplish.
IV-D2 Loss Function Experiment
Here we supply the results of RSEND trained with various combinations of losses. In Fig. 6(b), it shows the result of the combination of spatial consistency loss, exposure control loss, and color constancy loss, which is inspired by Zero-DCE [16]. However, the output is not lighted at all, showing that this combined loss is not suitable for our model. In Fig. 6(c), we use Charbonnier loss[31] to train our model and get this output, the loss is a smooth approximation of the L1 loss, to measure the pixel-wise difference between the enhanced image and the ground truth. The image is very close to our RSEND result in Fig. 6(e), but it is still not lighted enough with some loss of details in the darker regions. Based on Fig. 6(e), we can say that $\mathcal{L}_{vgg}$ plays its role by ensuring that the enhanced image maintains textural and structural similarity to natural images as perceived by the human visual system. Furthermore, we conducted an ablation study to evaluate the performance of all the loss functions combined in 6(d). We set the weights of $\mathcal{L}_{col+spa+exp}$ and $\mathcal{L}_{Charbonnier}$ to be 1 and $\mathcal{L}_{vgg}$ to be 0.5. This decision is based on the need to ensure that the perceptual quality of the enhanced images, as captured by $\mathcal{L}{vgg}$, is sufficiently emphasized without overwhelming the primary objectives of maintaining spatial consistency, proper exposure, and color constancy, as well as minimizing pixel-wise errors. By assigning a lower weight to $\mathcal{L}_{vgg}$, we aim to balance its influence against the other losses. We hoped that the combination would surpass the performance of any single loss function, but it appears that the other losses dragged down the performance of $\mathcal{L}_{vgg}$. This indicates that while $\mathcal{L}_{vgg}$ ensures the textural and structural similarity, the additional losses might have introduced conflicting optimization objectives, thereby diminishing the overall performance.
V Conclusion
@@
We propose an efficient and accurate CNN-based framework, RSEND, for low-light image enhancement and it can be trained end to end with paired images. RSEND leverages the power of Retinex theory and squeeze and excitation network to significantly enhance the representational power of the network without increasing much computing requirements. In RSEND, we make the model understand which parts are darker and require more attention by introducing the dark region detection module. After enhancing, we refine the output, which aims to fine-tune the details. After element-wise multiplication of reflectance and illumination, we add the original low-light image in the end, which serves as residual learning to maintain high similarity. Then we denoise the output image to ensure the final result is not only well-illuminated but also visually pleasing. For all the steps mentioned above, we utilize Squeeze and Excitation Network to better capture the details. The quantitative and qualitative experiments show that our RSEND outperforms all the CNN-based models (by 1.69 dB to 3.63 dB improvements) in PSNR and yields results that are close to or even better than the Transformer-based models (by 1.22 dB to 2.44 dB improvements) while using 3.93–9.78$\times$ fewer parameters compared to the previous state-of-the-art-works.